OpenStudy (anonymous):

Angle H and angle J are complementary. Angle H measures x + 5 and angle J measures 2x – 5. Find the measure of angle H.

OpenStudy (kobeni-chan):

Complementary angles add up to 90 degrees. Add Angle H and Angle J to equal 90. 3x = 90. Solve for x, and then substitute that value into the equation for angle H.

OpenStudy (anonymous):

90/3 = 30 2(30) -5= 60-5= 55